# Conformalized Late Fusion Multi-View Learning

### v1

@@ 4.2 MVCP Predict-Then-Optimize
where $p\in\mathbb{R}^{n}$ and $B>0$. The distributions $\mathcal{P}(C)$ and $\mathcal{P}(X\mid C)$ are again taken to be those from various simulation-based inference (SBI) benchmark tasks provided by [51], chosen as they have $\mathcal{P}(C\mid X)$ with complex structure, fully described in Appendix C. We specifically study the Gaussian Mixture distributional setup across increasing dimensions to highlight the capacity for MVCP to succeed even for complex combinations of distributions.
We then consider $K=2$ views, namely with $\mathcal{X}=\mathbb{R}^{d}$, $\mathcal{X}_{1}=\mathbb{R}^{1:d_{1}}$, and $\mathcal{X}_{2}=\mathbb{R}^{d_{1}+1:d}$. For each $\mathcal{X}_{k}$, we train a generative model $q_{k}(C\mid X_{k})$, specifically a normalizing flow. We then compare four distinct ways of formulating $\mathcal{C}(X)$, namely by conformalizing each view individually, by conformalizing a fused map, by fusing the scores via a sum, and by fusing with MVCP, all using the score given by Equation 3, specifically taking $g_{k}$ to be the $\ell^{2}$ metric. The “Pre-fused” strategy employed here is again a mean aggregation, i.e. a sample $\widehat{c}\sim\mathcal{F}(q_{1},q_{2})$ is drawn via $\widehat{c}_{1}\sim q_{1}(C\mid X_{1}),\widehat{c}_{2}\sim q_{2}(C\mid X_{2})$ and $\widehat{c}$ defined as $(\widehat{c}_{1}+\widehat{c}_{2})/2$.
As discussed in Section 3.3, the prediction regions are not explicitly constructed in regression settings, rendering the reporting of prediction region sizes impossible. We instead compare the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.2. This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting, with the upper bound seen as Equation 5 can trivially be solved by simply taking $w=0$. The results are presented in Table 2.
Table 2: Coverages across tasks for $\alpha=0.05$ are shown in the left table, where coverage was assessed over a batch of 500 i.i.d. test samples. Suboptimality gap proportions ($\Delta_{\%}$) are shown in the right table, with medians reported over a batch of 500 i.i.d. test samples and median average deviations in parentheses.
View 1 View 2 Pre-Fused Sum MVCP
Gaussian Mixture (1) 0.964 0.958 0.952 0.964 0.976
Gaussian Mixture (2) 0.938 0.938 0.93 0.95 0.94
Gaussian Mixture (3) 0.944 0.972 0.936 0.944 0.948
Gaussian Mixture (4) 0.942 0.94 0.914 0.95 0.924
View 1 View 2 Pre-Fused Sum MVCP
0.901 (0.1003) 0.886 (0.1144) 1.0 (0.0) 0.702 (0.1403) 0.542 (0.0946)
0.962 (0.0453) 1.0 (0.0155) 1.005 (0.0019) 0.893 (0.1084) 0.874 (0.0792)
1.011 (0.018) 0.997 (0.0208) 1.011 (0.0035) 0.838 (0.0735) 0.868 (0.047)
1.026 (0.0181) 0.987 (0.0327) 1.018 (0.0066) 1.003 (0.0144) 0.946 (0.0286)
We see that all methods produce the coverage guarantees in accordance to Theorem 3.1. The resulting suboptimality gaps additionally exhibit two notable characteristics. The first is that MVCP produces more informative prediction regions, and hence less conservative robust upper bounds, if the views in consideration are individually informative, that is, each has $\Delta_{\%}<1$. While the sum conformal fusion exhibits a similar pattern, its effective reduction is consistently less than that provided by MVCP. Additionally, in cases where one of the two views is uninformative, i.e. $\Delta_{\%}=1$, MVCP nearly retains the performance of the informative view, as the quantile envelope adapts in considering all projection directions. The sum fusion technique, on the other hand, significantly degrades from the performance of the optimal view.
4.3 Real Data Classification
We now conduct experiments on the Handwritten dataset [50]. This dataset consists of features of 2000 handwritten digits $(0-9)$ (200 examples per class) with 6 different views of each image extracted from a collection of Dutch utility maps, described fully in Appendix D. The baseline approaches parallel those presented in Section 4.1. The results are compiled in Table 3. As in the synthetic classification experiment, we see that considering multiple views produces significantly more informative prediction regions than solely conformalizing a single view, especially when done in the manner prescribed by MVCP.
Table 3: Coverages across tasks for $\alpha=0.10$ are shown in the left table, where coverage was assessed over a batch of 295 i.i.d. test samples. Average prediction set sizes are shown in the right table, averaged over a batch of 295 i.i.d. test samples with standard deviations in parentheses.
$K$ Best View Pre-Fused Sum Norm MVCP
2 0.90 0.91 0.91 0.91 0.90
3 0.90 0.90 0.90 0.90 0.90
4 0.91 0.90 0.90 0.90 0.90
5 0.91 0.91 0.91 0.90 0.90
6 0.90 0.91 0.90 0.90 0.91
Best View Pre-Fused Sum Norm MVCP
3.39 (0.09) 3.69 (0.11) 3.52 (0.13) 3.51 (0.13) 2.68 (0.17)
3.39 (0.09) 3.48 (0.09) 2.97 (0.09) 2.95 (0.09) 1.84 (0.10)
3.28 (0.14) 3.42 (0.05) 2.87 (0.05) 2.83 (0.05) 1.64 (0.12)
3.28 (0.14) 3.12 (0.07) 2.73 (0.08) 2.68 (0.07) 1.55 (0.06)
3.19 (0.17) 3.07 (0.11) 2.71 (0.06) 2.64 (0.05) 1.60 (0.07)
@@ 5 Discussion
We have presented MVCP, a framework for producing informative prediction regions in late-stage fusion pipelines. We additionally demonstrated the generality of the proposed pipeline in settings of both classification and robust predict-then-optimize regression tasks. This work suggests many directions for extension. One point of interest is the integration of MVCP with other downstream regression applications. For instance, [52] proposed an end-to-end differentiable extension to [26]: extending MVCP for integration to their framework would be of great interest. Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the optimization of MVCP to leverage it in settings of robust control.
6 Acknowledgements
We acknowledge the support of the College of Literature, Science, and the Arts at the University of Michigan via the Meet the Moment Research Initiative.
Appendix A Multivariate Visualization
We walk through a visual presentation of the approach below to supplement the textual description in the main text. As discussed, we start with a collection of multivariate calibration scores $\mathcal{S}_{C}$ with $s\in\mathcal{S}$ being $\in\mathbb{R}^{K}$. For the purposes of visualization in this section, we have $K=2$. We first partition the score evaluations $\mathcal{S_{C}}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$ into a subset used to define the pre-ordering and the remainder to define the multivariate quantile.
Figure 2: The calibration score evaluations are first split between those used to define the pre-ordering (green) and those used to define the final multivariate quantile (red).
We first wish to define the pre-ordering over $\mathcal{S}_{C}^{(1)}$. As described in the main text, the goal is to define this using an indexed family of sets $\mathcal{A}_{t}$ with index $t\in\mathbb{R}$, after which the multivariate quantile reduces to the univariate quantile formulation. To ensure the final envelope over $\mathcal{S}_{C}^{(2)}$ remains as tight as possible, we wish to define this family in a data-driven fashion. Critically, the shape of this tightest envelope around $\mathcal{S}_{C}^{(2)}$ will vary across $\alpha$, meaning we must define the family separately for each choice of $\alpha$. We expect the contour of the tightest $\alpha$ envelope for $\mathcal{S}_{C}^{(1)}$ will be similar to that over $\mathcal{S}_{C}^{(2)}$, motivating such a choice to define the indexing family. To do this, we project $\mathcal{S}_{C}^{(1)}$ along a number of directions, finding the $\beta$ quantile along each, in turn defining a half-plane, where $\beta$ is as described in Section 3.2.
@@ Appendix B Multivariate Score Coverage
Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable. Assume further that $K$ maps $s_{j}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined. Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\}_{(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}}$. For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$. Then, denoting $\mathcal{C}(X):=\{y:s(X,y)\in\mathcal{Q}(\alpha)\}$,
$\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha.$ (6)
@@
Appendix C Simulation-Based Inference Benchmarks
The benchmark tasks are a subset of those provided by [53]. For convenience, we provide brief descriptions of the tasks curated by this library; however, a more comprehensive description of these tasks can be found in their manuscript.
C.1 Gaussian Mixture
A mixture of two Gaussians, with one having a much broader covariance structure:
$\text{{Prior}: }\beta\sim\mathcal{U}(-10,10)$
$\text{{Simulator}: }x\mid w\sim 0.5\mathcal{N}(x\mid w,I)+0.5\mathcal{N}(x\mid w,.01I)$
For any Gaussian Mixture ($k$) distribution referenced in Section 4, we combined separate instances of this setup, where the typical setup has $\beta\in\mathbb{R}^{2}$. In other words, we generate $x:=[x_{1},...,x_{k}]$ by drawing a prior over a similarly partitioned $\theta:=[\theta_{1},...,\theta_{k}]$ and taking $x_{i}\sim\mathcal{P}(X_{i}\mid\theta_{i})$ according to the above distribution.
Appendix D Real Data Featurization
The data features of the Handwritten dataset [50] are as follow:
  • 47 Zernike moments
  • 76 Fourier coefficients of the character shapes
  • 6 morphological features
  • 240 pixel averages in $2\times 3$ windows
  • 64 Karhunen-Love coefficients
  • 216 profile correlations
Appendix E Training Details
All encoders were implemented in PyTorch [54] with a Neural Spline Flow architecture. The NSF was built using code from [55]. Specific architecture hyperparameter choices were taken to be the defaults from [55] and are available in the code. Optimization was done using Adam [56] with a learning rate of $10^{-3}$ over 5,000 training steps. Minibatches were drawn from the corresponding prior $\mathcal{P}(Y)$ and simulator $\mathcal{P}(X\mid Y)$ as specified per task in the preceding section. Training these models required between 10 minutes and two hours using an Nvidia RTX 2080 Ti GPUs for each of the SBI tasks.

### v2

@@ 4.3 CSA Predict-Then-Optimize
where $x\in\mathbb{R}^{T\times H\times W}$ are the previous precipitation readings, $w_{e}\in\mathbb{R}^{|\mathcal{E}|}$ the traffic proportion routed along road $e$, $c\in\mathbb{R}^{|\mathcal{E}|}$ the transit times anticipated across roads, $A\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{E}|}$ the graph incidence matrix, and $b\in\mathbb{R}^{|\mathcal{V}|}$ the vector that specifies the routing problem, in which $b_{s}=1,b_{t}=-1,$ and $b_{k}=0$ for $s$ the travel source node, $t$ the terminal node, and $k\notin\{s,t\}$ all other nodes.
We then consider two probabilistic models for traffic prediction, namely one based on the classical probabilistic Lagrangian integro-difference approach (STEPS) of [46] and one on the modern latent diffusion model (LDM) approach of [47]. As a result of the higher inference cost of the latter, we consider the setup where $J_{1}>J_{2}$, specifically with $J_{1}=4$ and $J_{2}=1$, which additionally serves to empirically highlight the flexibility of non-uniform sampling from predictors discussed in Section 3.2. As discussed in Section 4, the alternate aggregation strategies do not lend themselves for use in this setting. We, therefore, only compare CSA to the best model selection strategy, specifically by comparing to the separate conformalizations of the two constituent predictors. Each model was conformalized with Equation 3 taken to be the score, where $d_{k}$ was set to be the standard 2-norm across both predictors, such that $s_{1}(x,c)=\min_{j\in 1,...,J_{1}}||\widehat{c}_{1j}-c||_{2}$ for $q_{1}(C\mid x)$ and similarly for $q_{2}(C\mid x)$. We here evaluate the methods using the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.3. This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting.
Experiments were conducted with $|\mathcal{D}_{C}|=200$. The suboptimality was then computed across 100 i.i.d. test samples. To, therefore, assess the improvement offered by CSA, we conduct two paired t-tests, with hypotheses
$H_{0}:\Delta^{(\mathrm{CSA})}_{\%}=\Delta^{(\mathrm{STEPS})}_{\%}\quad H_{1}:\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$
and similarly for $\Delta^{(\mathrm{CSA})}_{\%}$ and $\Delta^{(\mathrm{LDM})}_{\%}$. The coverages and results of such t-tests are provided in Table 3, from which we find that CSA significantly reduces the suboptimality after accounting for Bonferroni multiple testing. We see that, while conformalization of either of the two views individually already produces the desired coverage, CSA produces more informative prediction regions, and hence less conservative robust upper bounds.
Table 3: Coverages across tasks for $\alpha=0.05$ are shown in the left table across both the individually conformalized and CSA fused approaches. P-values of the paired t-tests comparing gap proportions ($\Delta_{\%}$) are shown in the right table. Both were computed over 100 i.i.d. test samples.
STEPS LDM CSA
0.981 0.962 0.968
$H_{1}$ p-value
$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$ $3.61\times 10^{-4}$
$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{LDM})}_{\%}$ $9.50\times 10^{-4}$
@@ 5 Discussion
We have presented CSA, a framework for producing informative prediction regions in ensemble predictor pipelines. We additionally demonstrated the generality of the proposed pipeline in settings of both classification and robust predict-then-optimize regression tasks. This work suggests many directions for extension. One point of interest is the integration of CSA with other downstream regression applications. For instance, [48] proposed an end-to-end differentiable extension to [36]: extending CSA for integration to their framework would be of great interest. Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the use of CSA in settings of robust control.
Appendix A CSA Visual Walkthrough
We walk through a visual presentation of the approach below to supplement the textual description in the main text. We start with a collection of multivariate calibration scores $\mathcal{S}_{C}$ with $s\in\mathcal{S}$ being $\in\mathbb{R}^{K}$. For the purposes of visualization in this section, we have $K=2$. We first partition the score evaluations $\mathcal{S_{C}}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, with a subset $\mathcal{S}_{C}^{(1)}$ used to define the pre-ordering and the remainder $\mathcal{S}_{C}^{(2)}$ to define the multivariate quantile.
Figure 2: The calibration score evaluations are first split between those used to define the pre-ordering (green) $\mathcal{S}_{C}^{(1)}$ and those used to define the final multivariate quantile (red) $\mathcal{S}_{C}^{(2)}$.
We first wish to define the pre-ordering over $\mathcal{S}_{C}^{(1)}$. As described in the main text, the goal is to define this using an indexed family of sets $\mathcal{A}_{t}$ with index $t\in\mathbb{R}$, after which the multivariate quantile approach reduces to the univariate quantile formulation. To ensure the final envelope over $\mathcal{S}_{C}^{(2)}$ remains as tight as possible, we wish to define this family in a data-driven fashion. Critically, the shape of this tightest envelope around $\mathcal{S}_{C}^{(2)}$ will vary across $\alpha$, meaning we must define the family separately for each choice of $\alpha$. We expect the contour of the tightest $\alpha$ envelope for $\mathcal{S}_{C}^{(1)}$ will be similar to that over $\mathcal{S}_{C}^{(2)}$, motivating such a choice to define the indexing family. To do this, we project $\mathcal{S}_{C}^{(1)}$ along a number of directions, finding the $\beta$ quantile along each, in turn defining a half-plane, where $\beta$ is as described in Section 3.1.
@@ Appendix B Multivariate Score Coverage
Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable. Assume further that $K$ maps $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined. Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\mid(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}\}$. For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ induced by a collection of nested sets $\{\mathcal{A}_{t}\}_{t\geq 0}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$. Then, denoting $\mathcal{C}(x):=\{y:s(x,y)\in\mathcal{Q}(\alpha)\}$,
$\mathcal{P}_{X^{\prime},Y^{\prime}}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$ (6)
@@
Appendix C CSA Predict-Then-Optimize Visual Walkthrough
We now present a visual accompaniment of the predict-then-optimize algorithm presented in Section 3.2. We once again take $K=2$ for visual clarity in this walkthrough, where the predictors are as discussed in Section 3.2, namely assumed to be generative predictors $q_{k}(C\mid X)$ where the number of samples per predictor are fixed to be $\{J_{k}\}$. For illustration, we assume $J_{1}=5$ and $J_{2}=3$, meaning predictions with the first model are made by drawing 5 samples and 3 for the second. We assume the CSA calibration of Section 3.1 has already been performed, from which a collection of projection directions and quantiles $\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$ are available that implicitly define an acceptance region $\widehat{\mathcal{Q}}$. We further assume the individual predictor score functions are all the GPCP score given in Equation 3, with $d_{k}$ from Equation 3 specifically here taken to simply be the standard Euclidean 2-norm, giving
$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}||\widehat{c}_{kj}-c||.$ (10)
We now wish to compute $c^{*}=\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$. To do so, we must start by defining this region $\mathcal{C}(x)$ for the test point $x$, which we do by drawing the respective number of samples from the two models, producing samples $\{\widehat{c}_{1j}\}_{j=1}^{5}$ and $\{\widehat{c}_{2j}\}_{j=1}^{3}$, as shown in Figure 7.
Figure 7: Samples drawn from the two generative models $\{\widehat{c}_{1j}\}_{j=1}^{5}\sim q_{1}(C\mid x)$ (blue) and $\{\widehat{c}_{2j}\}_{j=1}^{3}\sim q_{2}(C\mid x)$ (green). Note that this is a visualization in the $\mathcal{C}$ space, i.e. not the space of multivariate scores.
By definition, $\forall c\in\mathcal{C}(x)$,
$u_{m}^{\top}\left(\min_{j_{1}=1,...,5}||\widehat{c}_{1j_{1}}-c||,\min_{j_{2}=1,...,3}||\widehat{c}_{2j_{2}}-c||\right)\leq\widehat{q}_{m}\qquad\forall m=1,...,M.$ (11)
As a result, we must have that, $\forall c\in\mathcal{C}(x)$, $\exists$ $j_{1}=1,...,5$ and $j_{2}=1,...,3$ such that $u_{m}^{\top}\left(||\widehat{c}_{1j_{1}}-c||,||\widehat{c}_{2j_{2}}-c||\right)\leq\widehat{q}_{m}$ $\forall m=1,...,M$. Solving for $c^{*}$, therefore, amounts to considering each pair $\vec{j}:=(j_{1},j_{2})\in\mathcal{J}$, where $\mathcal{J}:=\{1,...,5\}\times\{1,...,3\}$, and solving
$c_{\vec{j}}^{*}$ $:=\operatorname*{arg\,max}_{c}f(w,c)$ (12)
   s.t. $u_{m}^{\top}\left(||\widehat{c}_{1\vec{j}_{1}}-c||,||\widehat{c}_{2\vec{j}_{2}}-c||\right)\leq\widehat{q}_{m}\quad\forall m\in\{1,...,M\}$
Notice that, for any fixed $\vec{j}$, this is a standard convex optimization problem with a convex feasible region. We illustrate how the feasible region would be constructed for a fixed $\vec{j}$ in Figure 8. Note that the construction of this feasible is never explicitly done in practice and is only implicitly used by convex solver routines in practice.
Figure 8: A candidate $c\in\mathcal{C}$ is in prediction region if the projections of its distances $(d_{1},d_{2})$ from at least one pair of points indexed by $\vec{j}:=(j_{1},j_{2})$ is in the acceptance region $\widehat{\mathcal{Q}}$. Here, we illustrate a point $c$ that lies in the feasible region from its proximity to the $\vec{j}:=(4,1)$ pair of points. Note that the left is again a visualization over the $\mathcal{C}$ space, whereas the right is of the score space.
We can, therefore, then solve Equation 12 over all possible $\vec{j}\in\mathcal{J}$ and aggregate the maxima to compute $c^{*}$.
Appendix D Additional UCI Results
We consider those regression tasks from the UCI repository [45] that have at least 1,000 samples. The complete list of tasks considered herein is provided in Table 4.
Table 4: Tasks considered for empirical validation of prediction region efficiency and coverage, along with their relevant dimensions. All tasks are scalar regression tasks, i.e. $y\in\mathbb{R}$.
Dataset Observations Input Dimension
3droad 434874 3
airfoil 1503 5
bike 17379 17
parkinsons 5875 20
pol 15000 26
pumadyn32nm 8192 32
slice 53500 385
The complete collection of results is presented in Table 5. As in the main text, we see reduced conservatism of the CSA prediction regions compared to alternatives. We additionally notice that the single-stage CSA approach fails to retain the coverage guarantees.
Table 5: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d. test samples (10% of the dataset size). Standard deviations and means were computed across 10 randomizations of draws of the training, calibration, and test sets. Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, therefore, precluded from comparison.
Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
airfoil Coverage 0.95 (0.012) 0.981 (0.018) 0.954 (0.02) 0.969 (0.014) 0.962 (0.014) 0.939 (0.024) 0.969 (0.014) 0.854 (0.04) 0.952 (0.006)
Length 18.742 (0.313) 25.218 (1.05) 7.353 (0.672) 25.256 (1.424) 18.097 (0.565) 11.96 (0.634) 18.303 (0.787) 1.873 (0.21) 2.714 (0.141)
bike Coverage 0.963 (0.004) 0.959 (0.004) 0.948 (0.001) 0.955 (0.003) 0.955 (0.003) 0.932 (0.008) 0.959 (0.003) 0.946 (0.005) 0.955 (0.003)
Length 3.397 (0.131) 3.416 (0.123) 0.025 (0.0) 1.716 (0.104) 1.654 (0.11) 0.837 (0.055) 2.094 (0.09) 0.569 (0.097) 0.629 (0.105)
parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
Appendix E Robust Traffic Routing Setup
We replicate the experimental setup of [36], namely where a graph of Manhattan with corresponding nominal transit times was extracted using OSMnx [49]. Formally, the Manhattan graph is given as a tuple $(\mathcal{V},\mathcal{E})$, where the edge weights represent the transit times along the respective city roads. Such weights were assigned in a two-step process, namely by first making weather predictions and then using such weather predictions to then upweight the nominal transit times. In particular, precipitation forecasts were made from time-series observations of previous precipitations readings, specifically given over a map spatially resolved to $H\times W$ resolution. Precipitation forecasters, such as those considered in the experiments herein as given in [46] and [47], specifically map such previous observations to potential future trajectories. Formally, they define probabilistic models over some future time horizon $T_{f}$, from which probabilistic draws $\widetilde{Y}\in\mathbb{R}^{T_{f}\times H\times W}\sim\mathcal{P}(\widetilde{Y}\mid x)$ can be made, where $x\in\mathbb{R}^{T\times H\times W}$. Notably, we instead consider the probabilistic forecasts at some future fixed time point $T^{\prime}$, meaning the outcome of interest $Y\in\mathbb{R}^{H\times W}=\widetilde{Y}_{T^{\prime}}$
From a precipitation map, namely a spatially resolved reading $Y\in\mathbb{R}^{H\times W}$, we assign the final edge weights by first associating nodes to the closest pixel coordinate of the precipitation map. That is, denoting the pixel nearest to a vertex $v$ as $(p^{v}_{x},p^{v}_{y})$, the node is assigned the value at such a spatial location $Y_{p^{v}_{y},p^{v}_{x}}$. To, therefore, assign the edge weight, we average the weights of the edge endpoints and then weigh the nominal transit time. In particular, denoting the nominal transit time along such an edge $e$ between nodes $(s,t)$ as $\widetilde{c}_{e}$, the transit time with traffic was computed as
$c_{e}:=\widetilde{c}_{e}\cdot\exp\left\{\frac{Y_{p^{e_{s}}_{x},p^{v}_{y}}+Y_{p^{e_{t}}_{x},p^{e_{t}}_{y}}}{2}\right\}.$ (13)
Appendix F Conformal Aggregation Methods
We now describe the methods from [25] that were compared against experimentally, specifically the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches. As discussed in Section 2.4, these methods all follow the structural form of Equation 2 and largely differ in their choice of weights and thresholds. The standard majority-vote $\mathcal{C}^{M}$ is the most natural choice, defined by
$\mathcal{C}^{M}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>\frac{1}{2}\right\}.$ (14)
The randomized methods differ in that independent randomization is leveraged over the threshold, namely with:
$\mathcal{C}^{R}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>\frac{1}{2}+\frac{U}{2}\right\}$ (15)
$\mathcal{C}^{U}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>U\right\},$ (16)
for $U\sim\mathrm{Unif}([0,1])$. Notably, all these methods retain the guarantees typical of conformal prediction.